OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. x = 3y2

OpenStudy (anonymous):

Vertex: (0, 0); Focus:(1/2,0) ; Directrix: x = ; Focal width: 12 Vertex: (0, 0); Focus:(1/2,0) ; Directrix: x = ; Focal width: 0.33 Vertex: (0, 0); Focus:(0, 1/16) ; Directrix: x = ; Focal width: 0.33 Vertex: (0, 0); Focus:(1/16,0) ; Directrix: y = ; Focal width: 12

OpenStudy (anonymous):

somebody please help!